OpenStudy (anonymous):

prove f(x) has an inverse function and find the slope of the tangent line at point P on the graph of f^-1 f(x) = 4x^5 - (1/(x^3)), x>0; P(3,1)

OpenStudy (anonymous):

is this \[f(x)=4x^5-\frac{1}{x^3}\]?

OpenStudy (anonymous):

yes thats the correct equation :)

OpenStudy (anonymous):

ok do you know why it has an inverse?

OpenStudy (anonymous):

one to one, constantly increasing

OpenStudy (anonymous):

constantly increasing is not quite good enough, you have to also say it is continuous

OpenStudy (anonymous):

in fact it is not continuous, but it is on \(x>0\) so you need to include that fact as well

OpenStudy (anonymous):

sounds good, got that part down

OpenStudy (anonymous):

now finding the slope is easier than it looks. what you need to use is that \[(f^{-1})'=\frac{1}{f'(f^{-1}(x))}\]\]

OpenStudy (anonymous):

so you do not actually have to find the derivative of the inverse

OpenStudy (anonymous):

\((1,3)\) is on the graph of \(f\) and so \((3,1)\) is on the graph of \(f^{-1}\)

OpenStudy (anonymous):

so, \[f' = 20x^4 -(3/x^4) + c\] I am having a hard time finding the inverse thought

OpenStudy (anonymous):

a couple mistakes there

OpenStudy (anonymous):

but before we correct them, lets make sure one thing is clear you do not need to find the inverse function in fact, you probably cannot find it

OpenStudy (anonymous):

hmm okay, that would probably makes a lot of sense considering my struggles over finding it

OpenStudy (anonymous):

first off \[f'(x)=20x^4+\frac{3}{x^4}\] your sign was wrong also this is a derivative, not an integral, so there is no \(+C\) out at the end

OpenStudy (anonymous):

clear enough, right?

OpenStudy (anonymous):

oops, i read my paper wrong i had 20x^4 - -(3/x^4)

OpenStudy (anonymous):

ok good now look at the way to find the derivative of an inverse i wrote above, i will write it again \[(f^{-1})'(x)=\frac{1}{f'(f^{-1}(x))}\]

OpenStudy (anonymous):

but yes I understand, been a long day not sure whey I added the +C either :(

OpenStudy (anonymous):

and so of course \[(f^{-1})'(3)=\frac{1}{f'(f^{-1}(3))}\]

OpenStudy (anonymous):

we know everything we need to know to compute this, without finding an explicit expression for \(f^{-1}(x)\) in fact, you can do it in your head

OpenStudy (anonymous):

on the right hand side we need \(f^{-1}(3)\) usually you have to find that, but not in this case because they told you the point is \((3,1)\) so you know \(f^{-1}(3)=1\)

OpenStudy (anonymous):

so find \[1/ f'(1)\] ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

like i said, you can do it in your head \[f'(1)=20+3=23\] and so your answer is \(\frac{1}{23}\)

OpenStudy (anonymous):

Thank you for your help and you time, I really appreciate it!

OpenStudy (anonymous):

usually when people are given problems like this they spend 2 hours trying to find the inverse and get stuck, but notice how easy it is yw

OpenStudy (anonymous):

alot easier than i was expecting